### v1

# TexShape: Information Theoretic Sentence Embedding for Language Models

Abstract
"THIS PAPER IS ELIGIBLE FOR THE STUDENT PAPER AWARD." With the exponential growth in data volume and the emergence of data-intensive applications, particularly in the field of machine learning, concerns related to resource utilization, privacy, and fairness have become paramount. This paper focuses on the textual domain of data and addresses challenges regarding encoding sentences to their optimized representations through the lens of information-theory. In particular, we use empirical estimates of mutual information, using the Donsker-Varadhan definition of Kullback–Leibler divergence. Our approach leverages this estimation to train an information-theoretic sentence embedding, called TexShape, for (task-based) data compression or for filtering out sensitive information, enhancing privacy and fairness. In this study, we employ a benchmark language model for initial text representation, complemented by neural networks for information-theoretic compression and mutual information estimations. Our experiments demonstrate significant advancements in preserving maximal targeted information and minimal sensitive information over adverse compression ratios, in terms of predictive accuracy of downstream models that are trained using the compressed data.
Index Terms:
language model, sentence embedding, compression, privacy, fairness, mutual information.
I Introduction
The domain of machine learning (ML), particularly that of transformers and language models, has witnessed explosive growth over the past few years, e.g. [1, 2, 3] among many others. Specifically, the advancement of technologies such as ChatGPT [3], Bard [4], LLaMa [5], etc. have produced significant strides within the field of natural language processing. Essentially, the ability to extract representations from massive datasets and use such representations to classify and generate synthetic versions have found a multitude of applications, ranging from consumer engagement to enterprise productivity/efficiency [6]. This process of transforming massive datasets into (generative) models is very promising and powerful [7]. However, its connection with theoretical aspects of statistics and machine learning are currently poorly understood. In particular, we understand very little about the connections between information-theoretic quantities and concepts within (generative) ML.
Since its inception [8, 9], information theory has formed the basis on which all data processing systems, including communication, signal processing, and compression, have been benchmarked and characterized [10, 11]. Information theory has been used to understand both lossless and lossy compression of data as well as optimal mechanisms of (universal) compression of data sources. Indeed, methods to combine such information-theoretic understanding with representations in machine learning would be invaluable to improve the complexity and performance of ML systems.
This paper is structured to be a very early step in this direction for large language models (LLMs). In particular, our goal is to develop optimized and succinct text representations for ML through the use of information-theoretic tools. We built upon the encoding steps in LLMs, where the utilization of information theory can result in improved and compressed data representations. This combination forms a perfect union as information theory provides insights into compression and representation of data, while ML models generally desire compressed representations for improved performance at lower complexities. Ultimately, such a connection can potentially result in the design of an “optimal" encoder that has the best compressed representation for a desired level of performance. Furthermore, information theoretic tools can also be employed to succinctly and measurably capture privacy and fairness constraints within an ML setting.
Mutual information (MI) is an integral concept in information theory. In particular, MI is used to quantify the average related information between the outcomes of two random distributions [8], making it a great candidate for (task-oriented) sample representations [12]. However, one major drawback is that computing MI requires knowing the probability distributions that are often intractable to estimate in practice. To solve this problem, empirical estimation of MI has been heavily pursued with recent work attracting lots of attention[13, 14, 15, 16]. These estimations, although very promising, still face practical challenges with respect to stability, precision, and complexity. A framework, InfoShape [17], for these problems was recently presented to incorporate MI in designing privacy-preserving image representations. At a high level, our work expands upon InfoShape to present a nuanced methodology for utilizing MI estimation for semantic sentence embedding.
Our contributions can be summarized as follows:
  • We present a new information-theoretic architecture for semantic sentence embedding, based on bidirectional encoder representations from transformers (BERTs) [1], deep neural networks (DNNs), and variational MI inference [18, 19].
  • We utilize these information-theoretic embeddings to enable customized (task-based) text compression to enhance desired features within the overall ML system, such as better resource management, improved privacy, and/or greater fairness.
II Preliminaries and Problem Statement
In this paper, scalar are represented with lower-case letters, and matrices are represented with uppercase letters. Random variables are distinguished from deterministic ones using boldface style; for example $\bm{X}$ represents a random matrix.
II-A Preliminaries
Consider two random variables $\bm{X}\in\mathcal{X}$ and $\bm{X}\in\mathcal{Z}$, where $\mathcal{X}$ and $\mathcal{Z}$ indicate the space set of the two random variables. The MI value $\mathcal{I}(\bm{X};\bm{Z})$ quantifies how much information is obtained on average about $\bm{X}$ by observing $\bm{Z}$. By Shannon’s definition [8],
$\mathcal{I}(\bm{X};\bm{Z})=\sum_{X\in\mathcal{X}}\sum_{Z\in\mathcal{Z}}P_{\bm{X,Z}}(X,Z)\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)}),$
where $P_{\bm{X,Z}}(X,Z)$ is the joint probability mass function (PMF) of $\bm{X}$ and $\bm{Z}$, and $P_{\bm{X}}(X)$ and $P_{\bm{Z}}(Z)$ are PMFs for $\bm{X}$ and $\bm{Z}$, respectively. [^1] For continuous random variables, the summations are replaced with integrals, and PMFs are replaced with probability distribution functions (PDFs). The logarithm base determines the information unit, which is nat (natural base) in this paper.
An MI representation, derived from the Donsker-Varadhan definition of the Kullback-Leibler divergence [18], is given below,
$\mathcal{I}(\bm{X};\bm{Z}){=}{\sup_{F:\Omega\rightarrow\mathbb{R}}}E_{P_{\bm{X,Z}}}[F(\bm{X},\bm{Z})]{-}\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(\bm{X},\bm{Z})}]).$ (1)
Here, $\Omega=\mathcal{X}\times\mathcal{Z}$ is the product space of the two random variables and the sub-index of the expectation shows the distribution over which the expectation is taken. The supremum is taken over all functions such that the expectations are finite. Proof of this derivation appears in the appendix.
This optimization problem can be solved numerically using ML techniques, e.g., [13, 14, 15] among others. For this, the sufficiently rich set of functions over which the search is performed is modeled with a network parameterized with a set of weights. Then, the optimal parameter that reaches the (local) maximum can be identified via stochastic gradient descent (SGD) techniques [20]. Also, the two expectations are replaced with empirical averages over samples of a mini-batch that are drawn according to the joint distribution $P_{\bm{X,Z}}$ and the product of the marginal distributions $P_{\bm{X}}P_{\bm{Z}}$.
II-B Problem Description
Our problem is to design an encoder $T_{\Theta}:\mathcal{X}\rightarrow\mathcal{Z}$ such that when applied to each sentence $X\in\mathcal{X}$, the encoded sentence meets certain information theoretical properties. Here, $\Theta$ represents the trainable parameters of the encoder. The architecture of the encoder is fixed, and its trainable parameters are optimized as follows:
$\max_{\Theta}$ $\,\gamma\mathcal{I}(T_{\Theta}(\mathbf{X});\mathbf{X})+\sum_{i=1}^{N_{l}}\lambda_{i}\mathcal{I}(T_{\Theta}(\mathbf{X});L_{i}(\mathbf{X}))$ (2)
$-\sum_{j=1}^{N_{s}}\mu_{j}\mathcal{I}(T_{\Theta}(\mathbf{X});S_{j}(\mathbf{X})).$
The parameters $\{\gamma,\lambda_{1},\dots,\lambda_{N_{l}},\mu_{1},\dots,\mu_{N_{s}}\}$ take non-negative values, and are tuned to enable a desired balance between competing goals of the optimization problem.
Next, we describe the benefit of each term in the encoder optimization problem (2). The first term is dedicated to information theoretical compression to maximize the mutual information between a compressed sentence and its original representation. We note that an information-theoretical compression stands above the classical methods, where simple distortion measures are minimized (e.g., mean square error after decompression). The second term is dedicated to the utility by placing tunable attention on certain parts of the sentence content, making the embedding more representative of selected information. One benefit of such a scheme is in scenarios where losing information is inevitable, to preserve all critical information. For example, through task-based compression of pilot scripts, the script execution process in airplane applications can be expedited without compromising safety. The last term is related to filtering out sensitive information from the encoded embedding either to make sure this sensitive information is not used by unauthorized users when data is publicly published (privacy), or to make sure biases are not incorporated to models that are trained using encoded representations (fairness). For example, by employing task-based encoding on finance documents, one can facilitate the development of unbiased decision models for loan approval.
By appropriately setting the design parameters, a variety of interesting problems can be formed. For instance, when $\lambda_{i}=\mu_{j}=0\;\forall i,j$, the problem is reduced to a task-agnostic information-theoretical compression. When only $\lambda_{i}=0\;\forall i$, the problem is task-oriented and is basically filtering our sensitive information in a dataset. Finally, when $\gamma=0$, the task-oriented problem is a privacy-utility trade-off where sentences in a dataset are encoded so that they can be used to train a model to identify public information (utility purpose). However, they are dramatically less useful for training a model to identify some sensitive information (privacy purpose). In fact, the latter was discussed in [17] for image domain with one public label and one private label.
Figure 1: Architecture for designing and utilizing TexShape.
III Designing TexShape
In this section, we provide the architecture and the training procedure of TexShape: our novel information-theoretic sentence embedding framework. The overall design strategy is illustrated in Fig. 1, which includes two main components: the sentence encoder TexShape which is highlighted with blue borders, and an information-theoretic evaluator that is used as a differentiable loss function to train TexShape and is highlighted with red borders. The trained encoder can be utilized on any dataset coming from the distribution that is used in the training stage, to customized its information theoretic properties; for a variety of applications such as bandwidth-constrained transmission, private data-sharing, or unbiased model developments. The cost of using TexShape is equivalent to the cost of ML inference in NLP models.
Encoder. The encoder consists of an off-the-shelf sentence embedding which sequentially receives tokens of a sentence and outputs a vector that represent the whole sentence. There are a variety of choices for this non-trainable part of the encoder such as [1, 21]. The sentence embedding is then fed into a neural network that projects the original sentence embedding into lower-dimensional embedding with certain information theoretical properties. This neural network is the trainable part of the encoder that is designed using an information theoretical evaluator, described next.
Information-Theoretic Evaluator. This part consists of estimators for the MI terms that appear in the optimization problem (2). MI is defined between two random variables $\bm{A}$ and $\bm{B}$, e.g., between a sentence embedding and its label or two different embedding for a sentence. We follow [13, 15, 14] and use the Donsker-Varadhan formulation of MI to empirically estimate it using samples drawn from the joint distribution. This formulation states that MI can be obtained by finding a function that maximizes an expression that only depends on the first moment of probabilistic terms, see (1). In this direction, the first moments can be empirically determined with the mean of the samples, known as the law of large numbers. In addition, the set of functions over which the search is performed can be modeled with a neural network, which is proved to model any well-behaved function with appropriate weights; see the universal approximation theorem [22]. Finally, the optimization is performed using SGD, where at each step a batch of samples is used in a public dataset $D=\{(A_{i},B_{i})\}_{i\in|\mathcal{D}|}$ to estimate expectations and their gradients. Then, the maximum converged value indicates the MI estimation.
Training. To separate the iterations of training the sentence encoder and iterations of MI estimators, we call the earlier an ‘epoch’ and the latter an ‘iteration’. At each epoch, several MI values are estimated, each requiring many iterations. As verified by our simulation results, the number of iterations is notably lower for higher epochs compared to the initial epochs. The encoder is initialized randomly for the first epoch, and at each epoch, its weights are updated to improve its information theoretical properties. MI values and their gradients can be iteratively estimated using batches of a public dataset from the intended distribution. After several epochs, the encoder can be isolated to be used by data owners.
Use Cases. The encoded sentences can be used for a variety of task-agnostic or task-oriented applications. For example, when the encoder acts as a compression tool, the encoded data can be handled with less overhead in terms of storage, bandwidth, complexity, and latency, etc. When the encoder acts as an encryption tool, certain sensitive features are removed, and the shared encoded data cannot be used by an adversary to infer sensitive individual properties of each sample. Finally, when the encoder acts as a fairness tool, certain sensitive information can be removed from training data to ensure downstream models do not incorporate biases in their decision-making. In the next section, we provide results for each of these applications using various datasets.
IV Simulation Results
In this section, we conduct our experiments on three public text datasets to optimize their representations according to different forms of the objective we defined in (2).
Dataset A: Stanford Sentiment Treebank (GLUE/SST2) [23]. It comprises sentences extracted from $67{,}349$ movie reviews, each annotated with a binary “sentiment" label, being positive or negative, with $56{-}44$ distribution. We also generated a binary label per sample as “sentence-length" with $50{-}50$ distribution, which indicates whether a sample contains more than eight tokens or not. The training data was then partitioned into a $90{:}10$ split, for training and validation.
Dataset B: Corona-NLP [24]. It includes a set of Covid-related tweets, each annotated with a “sentiment" label that has five different possibilities, i.e., extremely negative, negative, neutral, positive, extremely positive. During a pre-processing, we quantized the sentiment label to be binary, indicating if the tweet is positive or negative, via excluding neutral samples and merging extremely negative with negative classes, and extremely positive with positive classes. As for the second label, we used the author’s geographical “location" and removed samples with unclear location information. Thus, we obtained per sample a binary label indicating if the author is from United States or not. After pre-processing, the training and validation dataset was reduced to $14{,}335$ and $1{,}364$ samples, where the distribution of location was $56{-}44$ for training and $50{-}50$ for validation, while the distribution of sentiment was $45{-}55$ for training $51{-}49$ for validation.
Dataset C: MultiNLI (Multi-Genre Natural Language Inference) [25]. This dataset includes premise-hypothesis pairs of sentences, labeled entailment, contradiction, or neutral, specifying their logical relationship. Moreover, each sample is labeled as government or telephone, based on the source and context of each sample. Following a minimal pre-processing, $160{,}698$ training samples and $3{,}911$ validation samples are obtained from this dataset. The distributions of the two labels are $33{-}33{-}$ 33 - 33 and $48{-}52$ for the training data; and $36{-}33{-}31$ and $50{-}50$ for the validation data, respectively.
We next describe details of the TexShape encoder introduced in Fig. 1. The original model for sentence embedding (the non-trainable part of the encoder) is a pre-trained MPNet model [21], which is built upon Bidirectional Encoder Representations from Transformers (BERT) [1]. The original sentence embedding gives a vector of size $768$ per sample. The trainable part is designed to be a DNN with $768$ input nodes, two hidden layers having $512$ and $256$ nodes in most cases, and an output layer that its cardinality is chosen based on the compression ratio, all utilizing the ReLU activation function. The training utilizes Adam optimizer with a learning rate $10^{-3}$, and has $20$ epochs. The loss function is based on the information theoretical evaluator.
For a network of MI estimation, we used a DNN with the number of input nodes being summation of the size of two random variables that their MI is estimated, and the number of output node being one. Each MI estimator has two intermediate layers with $64$ and $32$ nodes, respectively. The activation function for the intermediate nodes is ReLU. The learning rate to optimize the cluster of models that estimate MIs is $10^{-4}$, the optimizer is Adam, and the number of optimization iterations is $2{,}000$. The loss function used for training is described in (2), [^2] Certain regularization terms can be added for improved stability in the numerical optimization [14]. which act on the entire dataset to estimate the intended MIs. We ran our experiments with an NVIDIA RTX A5000 GPU with a memory of 24GB.
IV-A TexShape to Address Privacy-Utility Trade-Off
In this subsection, we utilize Dataset A and Dataset B. We aim to compress each datasets to filter out sensitive information, while preserving critical information that is needed for certain downstream tasks. Thus, we choose the objective function to be a simplified form of (2), where $\gamma=0$, $N_{l}=N_{s}=1$, $\lambda_{1}=1$ and $\mu_{1}=\mu$, resulting in,
$\max_{\Theta}\,\mathcal{I}(T_{\Theta}(\mathbf{X});L(\mathbf{X}))-\mu\mathcal{I}(T_{\Theta}(\mathbf{X});S(\mathbf{X})).$
Here, $L(\mathbf{X})$ denotes the “sentiment" label for Dataset A (SST-2) and “location" label for Dataset B (Corona-NLP), and $S(\mathbf{X})$ denotes the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. Each MI is computed between a compressed representation with size $128$ and a binary label, and thus the network they utilize has $129$ input nodes. Each MI estimation iteration employs a batch size equal to the length of the full data set, since larger batch sizes showed improved numerical stability.
Your Alt Text
Figure 2: 2D visualizations of sentences in Dataset A, colored based on their public label (left) and private label (right). The top panel is dedicated to the original embedding, and the bottom panel is dedicated to TexShape embedding.
Fig. 2 illustrates the T-SNE mapping obtained before and after employing the TexShape encoder trained with $\mu=0.3$ on Dataset A. The illustrations show that the task-oriented encoder makes distinguishing between samples based on their private label visually impossible, while preserving the visual separability between samples based on their public labels. The estimated MIs between the embedding and the public label is $0.5051$ and $0.4739$ for original and TexShape embeddings. On the other hand, the estimated MIs between the embedding and the private label is $0.5552$ and $0.1334$ for original and TexShape embeddings.
We also considered two classifiers that are trained to predict the public label and the private label of a compressed sample, respectively. We compared three different sentence embedding methods for Dataset B as the training data: Original (when the information-theoretic compression part of the encoder is bypassed), Random (when weights of the the trainable part of the encoder is chosen randomly), and TexShape with $\mu=0.4$. Fig. 3 depicts the predictive performance of these classifiers on unseen data, in terms of their ROC curves, for public label and private label, respectively. As for public label, the area under the ROC curve (AUROC) for TexShape even slightly improves upon the corresponding number for the original representation, suggesting that a compressed version can be more beneficial. As for private label, the AUROC for TexShape is around $0.59$, showing that this embedding can hardly be used by an adversary to identify sensitive information, even when similar public data exists. The performance of random embedding is worse than TexShape, showing lower predictive utility and higher leakage of private information.
Refer to caption Refer to caption
Figure 3: Validation ROC Curves for classifiers trained on Dataset B for two different prediction tasks.
IV-B TexShape for Information-Theoretic Compression
In this subsection, we evaluate the compression capability of TexShape without emphasis on or filtering any specific feature. We utilize Dataset C, and we aim to compress textual dataset by transforming the original embedding with size $768$ into a compressed embedding with lower dimensionality, while maximizing the mutual information between the original embedding and the encoded embedding. Thus, the objective function derived from (2) is simplified to $\max_{\Theta}\,\mathcal{I}(T_{\Theta}(\bm{X});\bm{X})$. During TexShape training, only the premise part of each sample was utilized.
For evaluation, we considered the generalization accuracy (ACC.) of two different classifiers, i.e., predictor of the logical relationship between a pair of sentences (Label 1), and predictor of the source of a sentence (Label 2). The classifiers are trained on various embedding of Dataset C and Table I summarizes the results, which manifests the power of our MI-based compression in preserving the predictive utility across these two tasks. We would like to note that training of the encoder was agnostic to any of these features, and regardless, TexShape was successful to preserve notably higher information regarding each of these two features, compared to an untrained (Random) compression.
TABLE I: Effect of information-theoretic compression on predictive utility of downstream classifiers.
Embedding Size Label 1 ACC. Label 2 ACC.
Original Embedding 768 69.3 99.4
Random Embedding 128 51.1 90.8
TexShape Embedding 128 61.0 95.3
Random Embedding 64 35.9 82.4
TexShape Embedding 64 57.3 95.5
IV-C TexShape for Compression with Fairness Consideration
In this subsection, we utilize Dataset A and Dataset B, and the focus is to keep maximal information per sample on average, while filtering out certain information that can cause unwanted biases when used for training classifiers. Thus, the objective function is derived from (2) is $\max_{\Theta}\mathcal{I}(T_{\Theta}(\mathbf{X});\mathbf{X})-5\mathcal{I}(T_{\Theta}(\mathbf{X});S(\mathbf{X}))$. Here, $S(X)$ represents the sensitive information in $X$ that needs to be removed, which is the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. We measure the MI between the embedding and the sensitive label as a measure of bias. Further, to evaluate the utility of the filtered dataset, we evaluated the performance of classifiers that are trained to predict “sentiment" label for Dataset A and “location" label for Dataset B. Table II summarizes the results: As we observe, the measure of bias is dramatically lower in the outputs of TexShape, compared to the original datasets. Such a benefit comes at the cost of a small decrease in the AUROC for a task that TexShape was agnostic to, compared to a similar setting where training was done using original dataset.
TABLE II: Measure of utility and bias before and after TexShape encoding.
Embedding Size Dataset A Dataset B
AUROC Bias AUROC Bias
Original 768 0.8342 0.5552 0.9566 0.5832
TexShape 128 0.7920 0.1939 0.9284 0.0711
V Conclusions
In this paper, we have integrated fundamental concepts from information theory and machine learning and proposed a semantic approach for sentence embedding in language models. Additionally, we have highlighted applications of our solution in areas such as lossy compression, privacy-preserving data sharing, and training unbiased models. Our work represents an initial step in understanding the relationship between information theory and the rapidly-growing field of large language models.
Acknowledgment
This research is supported in part by a grant from the CII (Construction Industry Institute) - OS2 program, MIT Portugal Program, and a grant by JMA Wireless.
Donsker-Varadhan Representation for MI
Let consider an arbitrary well-behaved function $F:\Omega\rightarrow\mathbb{R}$. We define another distribution $G:\Omega\rightarrow\mathbb{R}$ (called Boltzmann distribution in the literature) as follows:
$G(X,Z)=\frac{e^{F(X,Z)}}{E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}]}P_{\bm{X}}(X)P_{\bm{Z}}(Z).$
Because of linearity of expectation, we have,
$-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}]){=}\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)}){-F}(X,Z).$ (3)
Since the left hand side of (3) is deterministic, the right hand side must also be deterministic, thus the expectation does not change its value,
$-\log(E$ [${}_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])=$
$E_{P_{\bm{X,Z}}}[\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})]-E_{P_{\bm{X,Z}}}[F(X,Z)].$
Leading to,
$E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E$ [${}_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])=$
$E_{P_{\bm{X,Z}}}[\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})].$
On the other hand, by definition of mutual information, we have
$\mathcal{I}(\bm{X};\bm{Z})=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})].$ (4)
Therefore,
$\begin{split}\Delta&=\mathcal{I}(\bm{X};\bm{Z})-\left[E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])\right]\\ &=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})-\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})]\\ &=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{G(X,Z)})]\\ &=D_{\text{KL}}(P_{\bm{X,Z}}(X,Z)||G(X,Z))\geq 0\end{split}$
As a result,
$E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])\leq\mathcal{I}(\bm{X};\bm{Z}).$
The bound is achieved when $D_{\text{KL}}(P_{\bm{X,Z}}(X,Z)||G(X,Z))=0$. This happens only when $G(X,Z)=P_{\bm{X,Z}}(X,Z)$. Equivalently, when
$F(X,Z)=\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})+C,$ (5)
for a constant $C$.